# A Lightweight Video Anomaly Detection Model with Weak Supervision and Adaptive Instance Selection

### v1

@@ Abstract
Video anomaly detection is to determine whether there are any abnormal events, behaviors or objects in a given video, which enables effective and intelligent public safety management. As video anomaly labeling is both time-consuming and expensive, most existing works employ unsupervised or weakly supervised learning methods. This paper focuses on weakly supervised video anomaly detection, in which the training videos are labeled whether or not they contain any anomalies, but there is no information about which frames the anomalies are located. However, the uncertainty of weakly labeled data and the large model size prevent existing methods from wide deployment in real scenarios, especially the resource-limit situations such as edge-computing. In this paper, we develop a lightweight video anomaly detection model. On the one hand, we propose an adaptive instance selection strategy, which is based on the model’s current status to select confident instances, thereby mitigating the uncertainty of weakly labeled data and subsequently promoting the model’s performance. On the other hand, we design a lightweight multi-level temporal correlation attention module and an hourglass-shaped fully connected layer to construct the model, which can reduce the model parameters to only 0.56% of the existing methods (e.g. RTFM). Our extensive experiments on two public datasets UCF-Crime and ShanghaiTech show that our model can achieve comparable or even superior AUC score compared to the state-of-the-art methods, with a significantly reduced number of model parameters.
Index Terms:
Video Anomaly Detection; Weak Supervision; Adaptive Instance Selection; Lightweight Model.
I Introduction
@@
Video surveillance serves as a critical tool for identifying unexpected or abnormal events in many scenarios such as traffic monitoring and public safety management. Traditionally, video surveillance is heavily dependent on manual operations and of low intelligence [1, 2, 3]. For example, many cameras are installed in public venues such as stations and parks, to monitor unexpected or abnormal events, which generate huge amounts of videos. To check these videos manually is time-consuming and laborious. The rapid development of computer vision and deep learning technologies has spurred more and more research on video abnormal event detection or video anomaly detection (VAD), which enables the applications of automatic scene monitoring and intelligent early warning.
@@
Generally, there are three types of video anomaly detection methods, supervised, wealy-supervied and unsupervided [3]. Among them, weakly-supervised video anomaly detection (WSVAD) uses weakly-labelled data to train the model, where the training videos are labeled whether or not they contain any anomalies, but there is no information about which frames the anomalies are located. Thus, WSVAD inherits the advantages of supervised and unsupervised methods. On the one hand, WSVAD has better performance than unsupervised methods as some supervision is exploited. On the other hand, WSVAD is much cheaper and more efficient in acquiring the training data than supervised methods, because the latter requires to label each frame whether it contains anomalies. Therefore, WSVAD becomes a hot topic of video anomaly detection.
@@
WSVAD is typically based on Multiple Instance Learning (MIL) [3, 4]. Under the MIL framework, a video is viewed as a bag that consists of various clips, each of which is considered as an instance. For the training videos, the annotations are on the video level. That is, we know which videos have anomalies, but we do not know which clips (or instances) and frames have anomalies. In this context, WSVAD methods face two major challenges. The first challenge is the uncertainty of the weakly labeled data: we do not know both the number and the locations of anomalous clips in each anomalous video, which limits the full exploitation of the training anomalous data, thus resulting in unsatisfactory performance. The second challenge is the huge model size. The models of existing methods have too many parameters, making them difficult to be applied in resource-pressing scenarios, such as edge-computing applications. Existing methods have been mainly trying to tackle these two challenges.
@@
Figure 1: The framework of our method Light-WSVAD. Our model is based on the multi-instance learning (MIL) framework. Each video is divided into 32 consecutive clips (or instances), which are grouped into a positive instance bag (for abnormal videos) or a negative instance bag (for normal videos). Video features are extracted by I3D. A Multi-level Temporal correlation Attention (MTA) module is designed to capture time-related information, which is then input to a Hourglass-shaped Fully Connected layer (HFC) to calculate the score of each instance. The top-$K$ reliable instances are selected based on an Adaptive Instance Selection (AIS) strategy for subsequent loss calculation.
@@
Concretely, existing MIL based methods [3, 5, 4, 6] adopt a strategy to maintain balance between the number of anomalous videos and that of normal videos in the training set, and compute the loss by selecting the instance (or clip) of the highest anomaly score. This strategy aims to minimize the uncertainty of weakly labeled data. Typically, in real scenarios, the time duration of anomalous videos used for training usually exceeds 30 seconds, which suggests the presence of multiple anomalous clips. However, the selection of only the instance (clip) of the highest anomaly score might lead to under-utilization of all the available anomalous data. Statistical finding from RTFM [7] reveals that the mean number of anomalous clips within an anomalous video is approximately three. Therefore, choosing the three instances (clips) of the largest feature magnitude in both anomalous and normal videos for loss calculation can more effectively utilize the training data, thereby improving performance. However, our point of view is that different numbers of instances should be chosen for different datasets. Therefore, we propose an Adaptive Instance Selection (AIS) strategy, which selects the number of normal or anomalous instances adaptively based on the current status of the model in training, thereby it is able to accommodate different training datasets.
@@
On the other hand, there are relatively fewer lightweight models for VAD in the literature [8, 9], which usually compromise performance to embrace smaller models. In this paper, to devise a lightweight yet high-performance anomaly detection model, we develop a lightweight Multi-level Temporal correlation Attention (MTA) module. This module emphasizes the relationships between video instances (clips) of different time spans, thus making the model focus on the important instances in the videos. Furthermore, we design an Hourglass-shaped Fully Connected layer (HFC), which contains only half the parameters of a conventional fully connected layer (FC), yet does not degrade the model’s performance. Finally, Observing the limitation of the sparsity loss widely used in existing methods, we specifically introduce a more effective antagonistic loss.
@@
    We propose an adaptive sample selection strategy that alleviates the uncertainty problem of weakly-labelled data and improves model performance.
@@
The rest of this paper is organized as follows: Sec. II reviews the related works and highlights the differences of our work from the existing ones. Sec. III introduces the proposed method in detail. Sec. IV presents the results of performance evaluation, including performance comparison with existing works and ablation studies. Finally, Sec. V concludes the paper and pinpoints the future works.
II Related Work
@@
II-A Weakly Supervised Video Anomaly Detection
@@
Traditional anomaly detection methods typically assume that only normal training data is available and use hand-crafted features for one-class classification to solve the problem [10, 11, 12, 13]. With the development of deep learning, some unsupervised methods detect anomalies using features such as human pose and object detection learned by deep neural networks [14, 15, 16]. The essence of anomaly detection methods [14, 15, 16] based on generative models assumes that the model cannot reconstruct unseen abnormal videos/images well, and thus regards samples with high reconstruction errors as anomalies. However, due to the lack of prior knowledge about anomalies, these methods are prone to overfitting to training data and unable to distinguish normal and anomalous events.
@@
Some works [3, 5, 4, 6, 7, 17, 18, 19] have shown that utilizing partially labeled abnormal samples can achieve better performance than unsupervised methods. However, the cost of obtaining a large number of frame-level labels is prohibitively high. Therefore, some video anomaly detection methods apply video-level labels for weakly supervised training. Sultani et al. [3] proposed a method that uses video-level labels and introduced the large-scale weakly supervised video anomaly detection (WSVAD) dataset UCF-Crime. This makes WSVAD one of the mainstream research directions in video anomaly detection [4, 7, 20, 21].
@@
Existing Weakly supervised video anomaly detection methods are mainly based on multple-instance learning (MIL). Given that the training data possesses only video-level labels, these approaches typically employ the instances of the highest anomaly prediction score from both positive and negative bags for loss computation during training, leading to under-utilization of the training data. To address this issue, Zhong et al. [22] transformed WSVAD into a binary classification problem in the presence of label noise. They employed a Graph Convolutional Neural Network (GCN) [23] to eliminate label noise, thus enhancing data utilization and model performance. While this method improves model performance, the training computation cost associated with GCN and MIL is considerably high. Furthermore, it may make features unconstrained in the feature space, resulting in unstable performance. Furthermore, Tian et al. [7] amalgamated representation learning and anomaly score learning by devising Robust Temporal Feature Magnitude (RTFM) learning. They separately selected the three instances of the largest feature magnitude, in both abnormal and normal bags for loss calculation. Such an approach can effectively utilize the training data, thus achiveving better performance. However, this training data utilization strategy does not consider the diverse characteristics of different datasets.
@@
II-B Attention Network
Attention networks are initially used for machine translation [24] and later widely used for various computer vision tasks such as image classification [25], object detection [26], image segmentation [27], image captioning [28], and action recognition [29] etc., and have achieved excellent performance. Recently, attention networks have also been employed in weakly supervised tasks. Choe et al. [30] proposed an attention-oriented dropout layer, leveraging self-attention mechanisms to address the weakly supervised object localization problem. W-TALC [31] combines MIL and common activity similarity loss to train an attention module to solve the weakly supervised action localization problem in videos. Li et al. [32] improved the ability of weakly supervised anomaly detection models to extract relationships between video frames by using SE-attention [33]. However, this method, not specifically designed for video data, fails to enable the model to concentrate on the temporal correlation between consecutive instances (clips), and has a relatively large number of parameters.
@@ II-B Attention Network
III Methodology
This section presents the proposed Lightweight Weakly Supervised Video Anomaly Detection (Light-WSVAD in short) method. The framework of Light-WSVAD is shown in Fig. 1. Light-WSVAD mainly comprises a lightweight multi-level temporal correlation attention module, an hourglass-shaped fully connected layer, and an adaptive instance selection strategy to alleviate the uncertainty of weakly labeled data. In addition, we employ a more robust antagonistic loss to further optimize the model’s performance. In what follows, we first give a formal definition of the problem, then introduce the major modules of our method in detail.
III-A Problem Statement
In the context of weakly-supervised learning, anomaly detection can typically be regarded as a multiple-instance learning problem. Given a set of videos with video-level annotations (i.e., just labeling whether or not a video in the training dataset contains anomalous content), Weakly-supervised video anomaly detection (WSVAD) aims to train a model with the annotated data, which is able to predict whether there are anomalies in any new videos.
III-B Feature Extraction
@@
III-C Multi-level Temporal Correlation Attention
@@
$\begin{split}T_{attention}=(Conv1D[k],Conv1D[k-2],\dots,Conv1D[3])\\ \ast G(\chi)\end{split}$ (1)
In Equ. (1), $Conv1D[k]$ represents the convolution kernel function of size $k~{}(k\leq T)$, *** denotes the convolution operation, $\chi$ is the feature vector of a video and $G(\chi)$ is the feature vector after global pooling. In the one-dimensional convolution layer, the convolution operation between $Conv1D[k]$ and $G(\chi)$ is equivalent to sliding the convolution kernel, and multiplying it with the feature, then adding the convolution results. The calculation of $G(\chi)$ is as follows:
$G(\chi)=\frac{1}{WH}{\textstyle\sum_{i=1,j=1}^{W,H}}\chi_{ij}$ (2)
@@
III-D Hourglass-shaped Fully Connected Layer
@@
III-E Adaptive Instance Selection
@@
Step 3. We select the top-$K$ instances in terms of feature magnitude [7] from the positive and negative bags, respectively. Thus, the loss function is as follows:
@@
III-F Antagonistic Loss Function
Existing methods of weakly supervised anomaly detection predominantly employ smooth loss and sparsity loss for model optimization. Smooth loss assumes that there exists only slight discrepancy in features between consecutive instances, and thus the variation in anomaly scores is tiny, which conforms to the real scenario. On the other hand, the sparsity loss assumes that anomaly events in videos occur infrequently, therefore the mean anomaly score of instances in an anomalous video should approach zero. Although the assumption of sparsity is reasonable, it does not hold in actual model training. Owing to computational resource constraint, typically 32 consecutive instances are used to form a bag. In training, when anomalies are present in a bag, the proportion of abnormal instances cannot be overlooked. According to our preliminary analysis conducted on ShanghaiTech [22], the ratio of abnormal instances approximates 20%, and the mean score is around 0.2.
Based on the above analysis, in this paper, we design a new smooth loss as follows:

### v2

@@ Abstract
Video anomaly detection is to determine whether there are any abnormal events, behaviors or objects in a given video, which enables effective and intelligent public safety management. As video anomaly labeling is both time-consuming and expensive, most existing works employ unsupervised or weakly supervised learning methods. This paper focuses on weakly supervised video anomaly detection, in which the training videos are labeled whether or not they contain any anomalies, but lack information about the specific frames and quantities of anomalies. However, the uncertainty of weakly labeled data and the large model size prevent existing methods from wide deployment in real scenarios, especially the resource-limit situations such as edge-computing. In this paper, we develop a lightweight video anomaly detection model. On the one hand, we propose an adaptive instance selection strategy, which is based on the model’s current status to select confident instances, thereby mitigating the uncertainty of weakly labeled data and subsequently promoting the model’s performance. On the other hand, we design a lightweight multi-level temporal correlation attention module and an hourglass-shaped fully connected layer to construct the model, which can reduce the model parameters to only 0.56% of the existing methods (e.g. RTFM). Our extensive experiments on two public datasets UCF-Crime and ShanghaiTech show that our model can achieve comparable or even superior AUC score compared to the state-of-the-art methods, with a significantly reduced number of model parameters.
keywords:
[…]
{highlights}
A new Lightweight video anomaly detection model is proposed .
Weakly labeled data problem is mitigated by an adaptive sampling strategy .
A lightweight multi-level temporal correlation attention module is designed.
A lightweight hourglass-shaped fully connected layer is designed.
Extensive experiments have shown that the proposed method is both lightweight and effective.
1 Introduction
@@
surveillance serves as a critical tool for identifying unexpected or abnormal events in many scenarios such as traffic monitoring and public safety management. Traditionally, video surveillance is heavily dependent on manual operations and of low intelligence [1, 2, 3]. For example, many cameras are installed in public venues such as stations and parks, to monitor unexpected or abnormal events, which generate huge amounts of videos. To check these videos manually is time-consuming and laborious. The rapid development of computer vision and deep learning technologies has spurred more and more research on video abnormal event detection or video anomaly detection (VAD), which enables the applications of automatic scene monitoring and intelligent early warning.
@@
Generally, there are there types of video anomaly detection methods, supervided [4], unsupervised [5, 7, 6, 8, 9] and weakly-supervised [3, 10, 11, 12, 13, 14, 15]. Supervised video anomaly detection typically requires frame-level or even pixel-level labels, which incurs expensive training cost. Hence, there is a little related research in this direction. Unsupervised video anomaly detection uses unlabeled data to train models, with the lowest training cost, but exhibits poor performance. Weakly-supervised video anomaly detection (WVAD) uses weakly-labeled data to train the model, where the training videos are labeled whether or not they contain any anomalies, but there is no information about which frames the anomalies are located. Thus, WVAD inherits the advantages of supervised and unsupervised methods. On the one hand, WVAD has better performance than unsupervised methods as some supervision is exploited. On the other hand, WVAD is much cheaper and more efficient in acquiring the training data than supervised methods, because the latter requires to label each frame whether it contains anomalies. Therefore, WVAD becomes a hot topic of video anomaly detection.
@@
WVAD is typically based on Multiple Instance Learning (MIL) [3, 10]. Under the MIL framework, a video is viewed as a bag that consists of various clips, each of which is considered as an instance. For the training videos, the annotations are on the video level. That is, we know which videos have anomalies, but we do not know which clips (or instances) and frames have anomalies. In this context, WVAD methods face two major challenges. The first challenge is the uncertainty of the weakly labeled data: we do not know both the number and the locations of anomalous clips in each anomalous video, which limits the full exploitation of the training anomalous data, thus resulting in unsatisfactory performance. The second challenge is the huge model size. The models of existing methods have too many parameters, making them difficult to be applied in resource-pressing scenarios, such as edge-computing applications. Existing methods have been mainly trying to tackle these two challenges.
@@
Figure 1: The framework of our method Light-WVAD. Our model is based on the multi-instance learning (MIL) framework. Each video is divided into 32 consecutive clips (or instances), which are grouped into a positive instance bag (for abnormal videos) or a negative instance bag (for normal videos). Video features are extracted by I3D. A Multi-level Temporal correlation Attention (MTA) module is designed to capture time-related information, which is then input to a Hourglass-shaped Fully Connected layer (HFC) to calculate the score of each instance. The top-$K$ reliable instances are selected based on an Adaptive Instance Selection (AIS) strategy for subsequent loss calculation.
@@
Concretely, existing MIL based methods [3, 11, 10, 12, 16] adopt a strategy to maintain balance between the number of anomalous videos and that of normal videos in the training set, and compute the loss by selecting the instance (or clip) of the highest anomaly score. This strategy aims to minimize the uncertainty of weakly labeled data. Typically, in real scenarios, the time duration of anomalous videos used for training usually exceeds 30 seconds, which suggests the presence of multiple anomalous clips. However, the selection of only the instance (clip) of the highest anomaly score might lead to under-utilization of all the available anomalous data. Statistical finding from RTFM [13] reveals that the mean number of anomalous clips within an anomalous video is approximately three. Therefore, choosing the three instances (clips) of the largest feature magnitude in both anomalous and normal videos for loss calculation can more effectively utilize the training data, thereby improving performance. However, our point of view is that different numbers of instances should be chosen for different datasets. Therefore, we propose an Adaptive Instance Selection (AIS) strategy, which selects the number of normal or anomalous instances adaptively based on the current status of the model in training, thereby it is able to accommodate different training datasets.
@@
On the other hand, there are relatively fewer lightweight models for VAD in the literature [14, 15, 16], which usually compromise performance to embrace smaller models. In this paper, to devise a lightweight yet high-performance anomaly detection model, we develop a lightweight Multi-level Temporal correlation Attention (MTA) module. This module emphasizes the relationships between video instances (clips) of different time spans, thus making the model focus on the important instances in the videos. Furthermore, we design an Hourglass-shaped Fully Connected layer (HFC), which contains only half the parameters of a conventional fully connected layer (FC), yet does not degrade the model’s performance. Finally, Observing the limitation of the sparsity loss widely used in existing methods, we specifically introduce a more effective antagonistic loss.
@@
    We propose an adaptive sample selection strategy that alleviates the uncertainty problem of weakly-labeled data and improves model performance.
@@
The rest of this paper is organized as follows: Sec. 2 reviews the related works and highlights the differences of our work from the existing ones. Sec. 3 introduces the proposed method in detail. Sec. 4 presents the results of performance evaluation, including performance comparison with existing works and ablation studies. Finally, Sec. 5 concludes the paper and pinpoints the future works.
2 Related Work
@@
2.1 Weakly Supervised Video Anomaly Detection
@@
Traditional anomaly detection methods typically assume that only normal training data is available and use hand-crafted features for one-class classification to solve the problem [17, 18, 19, 20]. With the development of deep learning techniques, some unsupervised learning methods utilize deep neural networks to learn features such as human posture and optical flow [5, 21, 22, 23, 24], or utilize the difference in feature distribution between normal and abnormal samples [9] to perform abnormality detection. The essence of unsupervised video anomaly detection methods [5, 21, 22, 23, 24] lies in the assumption that anomalies are rare events, and therefore the model is trained to learn mainly the feature distribution of normal samples. Based on this assumption, the model then determines whether a test sample is an anomaly by calculating the reconstruction errors or feature distributions of the test sample and the normal samples. However, due to the lack of prior knowledge about anomalies, these methods are prone to overfitting to training data and unable to distinguish normal and anomalous events.
@@
Some works [3, 11, 10, 12, 13, 25, 26] have shown that utilizing partially labeled abnormal samples can achieve better performance than unsupervised methods. However, the cost of obtaining a large number of frame-level labels is prohibitively high. Therefore, some video anomaly detection methods apply video-level labels for weakly supervised training. Sultani et al. [3] proposed a method that uses video-level labels and introduced the large-scale weakly supervised video anomaly detection (WVAD) dataset UCF-Crime. This makes WVAD one of the mainstream research directions in video anomaly detection [10, 13, 27, 28].
@@
Existing Weakly supervised video anomaly detection methods are mainly based on multple-instance learning (MIL). Given that the training data possesses only video-level labels, these approaches typically employ the instances of the highest anomaly prediction score from both positive and negative bags for loss computation during training, leading to under-utilization of the training data. To address this issue, Zhong et al. [29] transformed WVAD into a binary classification problem in the presence of label noise. They employed a Graph Convolutional Neural Network (GCN) [30] to eliminate label noise, thus enhancing data utilization and model performance. While this method improves model performance, the training computation cost associated with GCN and MIL is considerably high. Furthermore, it may make features unconstrained in the feature space, resulting in unstable performance. Furthermore, Tian et al. [13] amalgamated representation learning and anomaly score learning by devising Robust Temporal Feature Magnitude (RTFM) learning. They separately selected the three instances of the largest feature magnitude, in both abnormal and normal bags for loss calculation. Such an approach can effectively utilize the training data, thus achiveving better performance. However, this training data utilization strategy does not consider the diverse characteristics of different datasets.
@@
2.2 Attention Network
Attention networks are initially used for machine translation [31] and later widely used for various computer vision tasks such as image classification [32], object detection [33], image segmentation [34], image captioning [35], and action recognition [36] etc., and have achieved excellent performance. Recently, attention networks have also been employed in weakly supervised tasks. Choe et al. [37] proposed an attention-oriented dropout layer, leveraging self-attention mechanisms to address the weakly supervised object localization problem. W-TALC [38] combines MIL and common activity similarity loss to train an attention module to solve the weakly supervised action localization problem in videos. Zhou et al. [6] proposed an attention that focuses on the foreground of an image to alleviate the foreground-background imbalance problem in anomaly detection. Li et al. [39] improved the ability of weakly supervised anomaly detection models to extract relationships between video frames by using SE-attention [40]. However, this method, not specifically designed for video data, fails to enable the model to concentrate on the temporal correlation between consecutive instances (clips), and has a relatively large number of parameters.
@@ 2.2 Attention Network
3 Methodology
This section presents the proposed Lightweight Weakly Supervised Video Anomaly Detection (Light-WVAD in short) method. The framework of Light-WVAD is shown in Fig. 1. Light-WVAD mainly comprises a lightweight multi-level temporal correlation attention module, an hourglass-shaped fully connected layer, and an adaptive instance selection strategy to alleviate the uncertainty of weakly labeled data. In addition, we employ a more robust antagonistic loss to further optimize the model’s performance. In what follows, we first give a formal definition of the problem, then introduce the major modules of our method in detail.
3.1 Problem Statement
In the context of weakly-supervised learning, anomaly detection can typically be regarded as a multiple-instance learning problem. Given a set of videos with video-level annotations (i.e., just labeling whether or not a video in the training dataset contains anomalous content), Weakly-supervised video anomaly detection (WVAD) aims to train a model with the annotated data, which is able to predict whether there are anomalies in any new videos.
3.2 Feature Extraction
@@
3.3 Multi-level Temporal Correlation Attention
@@
$\begin{split}T_{attention}=(Conv1D[k],Conv1D[k-2],\dots,\\ Conv1D[3])\ast G(\chi)\end{split}$ (1)
In Equ. (1), $Conv1D[k]$ represents the convolution kernel function of size $k~{}(k\leq T)$, * denotes the convolution operation, $\chi$ is the feature vector of a video and $G(\chi)$ is the feature vector after global pooling. In the one-dimensional convolution layer, the convolution operation between $Conv1D[k]$ and $G(\chi)$ is equivalent to sliding the convolution kernel, and multiplying it with the feature, then adding the convolution results. The calculation of $G(\chi)$ is as follows:
$\begin{split}G(\chi)=\frac{1}{WH}{\textstyle\sum_{i=1,j=1}^{W,H}}\chi_{ij}\end{split}$ (2)
@@
3.4 Hourglass-shaped Fully Connected Layer
@@
3.5 Adaptive Instance Selection
@@
Step 3.The feature magnitudes [13] of the top-$K$ instances with the highest scores are selected separately from the positive and negative sets, and corresponding optimizations are performed accordingly. Thus, the loss function is as follows:
@@
3.6 Antagonistic Loss Function
Existing methods of weakly supervised anomaly detection predominantly employ smooth loss and sparsity loss for model optimization. Smooth loss assumes that there exists only slight discrepancy in features between consecutive instances, and thus the variation in anomaly scores is tiny, which conforms to the real scenario. On the other hand, the sparsity loss assumes that anomaly events in videos occur infrequently, therefore the mean anomaly score of instances in an anomalous video should approach zero. Although the assumption of sparsity is reasonable, it does not hold in actual model training. Owing to computational resource constraint, typically 32 consecutive instances are used to form a bag. In training, when anomalies are present in a bag, the proportion of abnormal instances cannot be overlooked. According to our preliminary analysis conducted on ShanghaiTech [29], the ratio of abnormal instances approximates 20%, and the mean score is around 0.2. Based on the above analysis, the use of smoothing loss is retained in this study. The implementation of smoothing loss is defined by the following formula: